OpenStudy (anonymous):

which of the following is a polynomial with roots -√3, √3, and -2 a) x^3-2x^2-3x+6 b) x^3-3x^2-5x+15 c) x^3+2x^2-3x-6 d)x^3+3x^2-5x-15

OpenStudy (anonymous):

@bibby always coming to the rescue lol

OpenStudy (bibby):

heh

OpenStudy (bibby):

\((x-(-2))*(x--(\sqrt3))(x-\sqrt3)=(x+2)(x+\sqrt3)(x-\sqrt3)\)

OpenStudy (bibby):

I'd say use something like wolfram alpha to compute that instead of multiplying it out but I just realized multiplying it out isn't that hard wolfram: http://www.wolframalpha.com/input/?i=(x%2B2)(x%2Bsqrt3)(x-sqrt3)&t=crmtb01 \(x^2+2x-3x-6\)

OpenStudy (bibby):

or we can do \((x+2)(x+\sqrt3)(x-\sqrt3)(x+2)(x^2\cancel{+\sqrt3-\sqrt3}-3)\\=(x+2)(x^2-3)\)

OpenStudy (bibby):

foiled out